# TabSTAR: A Foundation Tabular Model With Semantically Target-Aware Representations

### v1

@@ F.1 Dataset Level Performance
As discussed in §5, TabPFN-v2 is unable to run over 4 datasets: C12, because it is a multiclass problem with more than 10 classes, and C02, C14 and R01 because they support inference for up to 500,000 cells. Attempts to run the model over a subset of the examples led to a significantly worse performance, and thus we decide not to report them to allow a fair comparison. Furthermore, CARTE is unable to run over 15 of the datasets in the benchmark due a known bug [^37] https://github.com/soda-inria/carte/issues/23 in their implementation for the PowerTransformation, which struggles in the presence of features with too less unique values.
Table 17: Classification performance per dataset (up to 10K). The top performance score is bolded first, and then all scores are rounded. We report average AUROC with 95% CIs.
ID CARTE CatB CatB-T RF TabPFN TabSTAR XGB XGB-T
C01 88.4±0.3 90.2±0.3 90.3±0.4 88.8±0.3 90.3±0.3 90.8±0.3 89.3±0.4 90.2±0.4
C02 97.3±0.5 97.4±0.4 96.3±0.5 97.9±0.5 97.2±0.3 97.6±0.3
C03 82.5±0.3 82.0±0.3 81.2±0.6 77.3±0.3 82.4±0.3 83.0±0.3 80.5±0.3 82.1±0.3
C04 84.5±1.9 85.4±1.8 82.4±3.0 88.3±1.5 83.7±2.3 80.7±2.9 85.4±2.0
C05 88.5±0.9 90.6±1.0 90.9±0.5 88.0±1.1 91.2±0.7 91.3±0.8 88.1±1.1 90.3±0.6
C06 80.9±1.5 81.3±1.3 82.6±1.1 73.6±1.2 87.7±0.5 87.0±0.6 81.5±1.0 83.7±0.9
C07 90.1±0.4 90.3±0.3 90.6±0.3 85.1±0.7 89.8±0.4 91.5±0.3 87.2±0.6 89.3±0.5
C08 90.8±1.6 93.0±0.9 93.2±1.0 90.2±1.3 91.3±1.1 93.5±1.5 91.9±1.1 94.4±0.7
C09 99.8±0.1 99.8±0.1 99.6±0.2 99.8±0.1 99.3±0.1 99.7±0.1 99.8±0.1
C10 86.7±0.2 87.0±0.2 84.6±0.2 87.6±0.4 89.1±0.2 85.4±0.3 86.8±0.2
C11 78.4±0.5 79.7±0.5 80.7±0.4 76.2±0.6 81.4±0.4 79.2±0.5 78.2±0.6 80.2±0.5
C12 96.0±0.2 97.5±0.1 97.7±0.1 95.3±0.2 98.3±0.1 96.6±0.2 97.3±0.1
C13 70.9±0.8 73.7±1.1 74.1±1.0 71.1±1.1 72.3±0.9 75.0±0.7 70.2±0.8 74.1±1.1
C14 83.5±0.3 84.0±0.5 80.1±0.3 84.0±0.3 81.4±0.4 83.2±0.4
Table 18: Classification performance per dataset (above 10K). The top performance score is bolded first, and then all scores are rounded. We report average AUROC with 95% CIs.
ID CatB-T CatB-T-U TabPFN TabSTAR TabSTAR-U XGB-T XGB-T-U
C01 90.3±0.4 90.5±0.4 90.3±0.3 90.8±0.3 91.2±0.3 90.2±0.4 90.4±0.3
C02 97.4±0.4 98.3±0.3 97.9±0.5 98.4±0.2 97.6±0.3 98.2±0.3
C03 81.2±0.6 81.5±0.8 82.4±0.3 83.0±0.3 83.8±0.5 82.1±0.3 82.3±0.3
C07 90.6±0.3 91.0±0.4 89.8±0.4 91.5±0.3 91.9±0.3 89.3±0.5 89.9±0.5
C08 93.2±1.0 93.1±1.2 91.3±1.1 93.5±1.5 95.1±0.9 94.4±0.7 94.4±0.9
C09 99.8±0.1 99.8±0.1 99.8±0.1 99.3±0.1 99.5±0.1 99.8±0.1 99.8±0.1
C11 80.7±0.4 81.8±0.4 81.4±0.4 79.2±0.5 81.0±0.5 80.2±0.5 81.4±0.4
C12 97.7±0.1 98.5±0.1 98.3±0.1 99.1±0.1 97.3±0.1 98.4±0.1
C13 74.1±1.0 76.7±0.8 72.3±0.9 75.0±0.7 77.9±1.0 74.1±1.1 76.9±0.9
C14 84.0±0.5 84.8±0.4 84.0±0.3 85.1±0.3 83.2±0.4 84.2±0.4
Table 19: Regression performance per dataset. The top performance score is bolded first, and then all scores are rounded. We report average $R^{2}$ with 95% CIs.
ID CARTE CatB CatB-T RF TabPFN TabSTAR XGB XGB-T
R01 100±0.0 100±0.0 100±0.0 100±0.0 100±0.0 99.9±0.1 100±0.0
R02 98.3±1.0 98.3±0.9 98.0±1.0 98.4±0.9 98.0±1.1 97.9±1.1 98.3±0.9
R03 71.8±0.5 73.8±0.4 74.1±0.3 68.4±0.6 74.9±0.4 70.9±0.8 69.1±0.5 73.7±0.5
R04 86.2±7.4 85.1±7.1 85.1±6.8 86.1±5.2 81.5±8.7 85.8±7.5 87.5±3.7
R05 93.2±0.8 89.5±1.3 90.0±1.2 86.7±1.3 92.7±0.9 93.2±1.2 87.1±1.4 90.1±1.3
R06 97.7±0.8 98.1±0.5 98.2±0.5 97.2±1.0 98.5±0.6 98.1±0.3 97.1±0.7 97.9±0.7
R07 71.6±1.3 66.9±0.8 67.8±0.8 61.5±1.2 62.2±1.3 71.1±1.8 61.9±1.5 68.8±0.8
R08 93.2±0.8 93.0±0.7 93.2±0.7 92.2±0.7 93.9±0.7 92.8±0.6 92.3±0.7 93.1±0.7
R09 89.2±0.5 89.2±0.6 89.3±0.4 88.6±0.7 89.8±0.5 88.8±0.5 88.2±0.7 89.3±0.5
R10 99.5±0.2 99.0±0.6 98.9±0.6 98.1±0.3 99.1±0.3 99.5±0.1 98.6±0.3 99.4±0.2
R11 94.2±0.9 94.0±0.9 92.5±1.3 94.5±0.9 93.9±1.1 93.3±1.1 94.2±0.8
R12 98.5±0.3 98.5±0.2 98.0±0.3 98.5±0.2 98.3±0.2 98.2±0.3 98.5±0.2
R13 52.8±2.1 57.6±2.7 58.2±2.6 51.8±3.5 53.8±2.7 53.4±3.0 49.0±3.3 57.5±2.7
R14 96.6±0.2 97.3±0.1 97.3±0.1 96.8±0.1 96.2±0.2 96.4±0.1 97.1±0.1 97.1±0.1
R15 79.5±0.9 79.9±1.1 77.3±1.2 79.4±1.0 78.9±1.5 76.8±1.2 80.3±1.0
R16 98.7±0.0 98.7±0.0 98.7±0.0 97.7±0.1 98.8±0.0 98.7±0.0 97.7±0.1 97.8±0.1
R17 95.3±2.0 94.8±2.4 95.2±2.0 94.2±2.5 95.1±2.1 94.7±2.3 94.5±2.6 95.4±1.9
R18 98.1±0.4 98.2±0.4 98.2±0.4 97.7±0.4 93.1±1.3 97.4±0.6 97.5±0.8 98.3±0.4
R19 85.5±0.8 85.8±0.9 85.3±0.9 84.0±1.1 83.8±1.1 83.9±1.0 82.4±2.0
R20 96.4±0.5 96.0±0.6 96.2±0.5 95.7±0.6 93.9±1.0 95.6±0.9 95.3±0.9 96.2±0.5
R21 92.3±1.2 92.5±1.1 92.6±1.1 91.8±1.1 93.3±1.1 92.3±1.0 91.6±1.0 92.6±1.0
R22 45.2±5.5 45.2±5.9 39.1±6.4 43.8±5.5 39.7±5.5 36.1±7.1 46.8±5.8
R23 85.0±1.8 85.4±1.2 85.5±1.1 81.8±1.5 84.8±1.6 83.0±1.6 83.3±1.2 85.1±1.2
R24 86.0±0.6 84.0±0.8 85.8±0.7 79.7±1.0 70.3±1.6 81.8±1.8 82.9±0.7 85.9±0.9
R25 94.3±0.6 95.4±0.5 95.3±0.5 94.8±0.5 86.7±1.0 94.8±0.5 94.6±0.6 95.4±0.5
R26 99.8±0.1 99.4±0.1 99.4±0.1 99.0±0.2 99.8±0.1 99.6±0.1 99.1±0.2 99.5±0.1
R27 81.9±1.6 85.2±1.3 85.3±1.4 84.8±1.3 82.2±1.6 82.0±1.6 83.2±1.2 85.5±1.2
R28 52.5±2.6 53.4±2.7 53.3±2.8 46.0±2.7 61.7±2.1 51.5±2.9 45.3±2.3 53.5±2.7
[…]

### v2

@@ F.1 Technical Limitations
As discussed in §5, TabPFN-v2 is unable to run on 4 datasets: C12, because it is a multiclass problem with more than 10 classes, and C02, C14 and R01 because they support inference for up to 500,000 cells. Attempts to run the model over a subset of the examples led to a significantly worse performance, and thus we decide not to report them to allow a fair comparison.
Additionally, CARTE is unable to run over 15 of the datasets in the benchmark due to a known bug [^41] https://github.com/soda-inria/carte/issues/23 in their implementation for the PowerTransformation, which struggles in the presence of features with too less unique values. Furthermore, TabDPT’s performance over a few regression datasets is especially low. Since we focus on classification tasks, we exclude this model’s performance on this task. Finally, RealMLP, fails to run on R01 in the unlimited setting due to memory constraints.
F.2 Dataset Level Performance
We report AUROC for classification and $R^{2}$ for regression, with 95% CIs computed over the 10 runs for each dataset. Tables  19 and 20 summarize classification performance on datasets with up to 10K and over 10K examples, respectively. Tables LABEL:tab:reg_10k_dataset_performance and 22 to the same for regression tasks. For conciseness, datasets are referred by their ID from Appendix D.
Table 19: Classification performance per dataset (up to 10K). Best score before rounding is bolded. We report average AUROC with 95% CIs. Models: CARTE (CRT), CatBoost (CTB), TabICL (ICL), LightGBM (LGB), RealMLP (MLP), TabDPT (DPT), TabM (TBM), TabPFNv2 (PFN), RandomForest (RF), TabSTAR (STR), XGBoost (XGB). Tuned models are marked with a ’+’.
ID CRT CTB CTB+ DPT ICL LGB LGB+ MLP MLP+ PFN RF STR TBM TBM+ XGB XGB+
C01 88.4 $\pm$0.3 90.2 $\pm$0.3 90.3 $\pm$0.4 90.2 $\pm$0.3 90.7 $\pm$0.4 89.5 $\pm$0.4 89.6 $\pm$0.4 90.2 $\pm$0.5 90.2 $\pm$0.3 90.3 $\pm$0.3 88.8 $\pm$0.3 90.8 $\pm$0.3 90.5 $\pm$0.3 90.5 $\pm$0.3 89.3 $\pm$0.4 90.2 $\pm$0.4
C02 97.3 $\pm$0.5 97.4 $\pm$0.4 92.3 $\pm$0.6 95.2 $\pm$0.4 97.2 $\pm$0.5 97.3 $\pm$0.5 96.4 $\pm$0.5 96.6 $\pm$0.5 96.3 $\pm$0.5 97.9 $\pm$0.5 96.8 $\pm$0.5 97.2 $\pm$0.4 97.2 $\pm$0.3 97.6 $\pm$0.3
C03 82.5 $\pm$0.3 82.0 $\pm$0.3 81.2 $\pm$0.6 85.4 $\pm$0.3 81.4 $\pm$0.3 79.8 $\pm$0.3 79.9 $\pm$0.3 82.6 $\pm$0.4 82.7 $\pm$0.3 82.4 $\pm$0.3 77.3 $\pm$0.3 83.0 $\pm$0.3 83.3 $\pm$0.3 83.8 $\pm$0.3 80.5 $\pm$0.3 82.1 $\pm$0.3
C04 84.5 $\pm$1.9 85.4 $\pm$1.8 83.3 $\pm$2.8 86.9 $\pm$2.3 83.0 $\pm$2.3 82.8 $\pm$2.1 82.4 $\pm$3.1 82.8 $\pm$2.2 88.3 $\pm$1.5 82.4 $\pm$3.0 83.7 $\pm$2.3 85.0 $\pm$3.1 85.1 $\pm$2.6 80.7 $\pm$2.9 85.4 $\pm$2.0
C05 88.5 $\pm$0.9 90.6 $\pm$1.0 90.9 $\pm$0.5 92.2 $\pm$0.6 92.9 $\pm$0.5 88.2 $\pm$1.3 88.4 $\pm$0.6 91.0 $\pm$0.8 91.4 $\pm$0.6 91.2 $\pm$0.7 88.0 $\pm$1.1 91.3 $\pm$0.8 91.1 $\pm$0.7 90.7 $\pm$0.6 88.1 $\pm$1.1 90.3 $\pm$0.6
C06 80.9 $\pm$1.5 81.3 $\pm$1.3 82.6 $\pm$1.1 73.5 $\pm$2.2 74.3 $\pm$1.8 81.9 $\pm$1.1 82.8 $\pm$0.9 75.6 $\pm$1.4 77.5 $\pm$1.4 87.7 $\pm$0.5 73.6 $\pm$1.2 87.0 $\pm$0.6 79.0 $\pm$1.7 79.2 $\pm$1.5 81.5 $\pm$1.0 83.7 $\pm$0.9
C07 90.1 $\pm$0.4 90.3 $\pm$0.3 90.6 $\pm$0.3 89.9 $\pm$0.2 86.7 $\pm$0.3 88.9 $\pm$0.4 89.3 $\pm$0.4 90.0 $\pm$0.3 90.2 $\pm$0.2 89.8 $\pm$0.4 85.1 $\pm$0.7 91.5 $\pm$0.3 91.1 $\pm$0.3 91.1 $\pm$0.3 87.2 $\pm$0.6 89.3 $\pm$0.5
C08 90.8 $\pm$1.6 93.0 $\pm$0.9 93.2 $\pm$1.0 93.2 $\pm$1.4 95.1 $\pm$0.7 93.1 $\pm$0.7 93.2 $\pm$0.9 92.1 $\pm$0.9 91.9 $\pm$1.3 91.3 $\pm$1.1 90.2 $\pm$1.3 93.5 $\pm$1.5 92.1 $\pm$1.3 92.1 $\pm$1.5 91.9 $\pm$1.1 94.4 $\pm$0.7
C09 82.4 $\pm$1.3 82.6 $\pm$1.4 81.3 $\pm$1.2 84.3 $\pm$1.0 81.1 $\pm$1.2 82.9 $\pm$1.3 82.6 $\pm$1.4 81.3 $\pm$2.5 82.5 $\pm$1.1 76.4 $\pm$1.4 94.0 $\pm$0.8 81.8 $\pm$4.1 83.6 $\pm$1.0 79.4 $\pm$1.6 83.2 $\pm$1.2
C10 86.7 $\pm$0.2 87.0 $\pm$0.2 87.2 $\pm$0.2 87.7 $\pm$0.2 85.7 $\pm$0.3 85.9 $\pm$0.3 87.7 $\pm$0.2 87.6 $\pm$0.2 87.6 $\pm$0.4 84.6 $\pm$0.2 89.1 $\pm$0.2 87.5 $\pm$0.3 87.5 $\pm$0.3 85.4 $\pm$0.3 86.8 $\pm$0.2
C11 78.4 $\pm$0.5 79.7 $\pm$0.5 80.7 $\pm$0.4 79.6 $\pm$0.4 81.5 $\pm$0.5 78.9 $\pm$0.5 79.2 $\pm$0.4 80.4 $\pm$0.4 79.6 $\pm$0.5 81.4 $\pm$0.4 76.2 $\pm$0.6 79.2 $\pm$0.5 81.2 $\pm$0.3 81.5 $\pm$0.3 78.2 $\pm$0.6 80.2 $\pm$0.5
C12 96.0 $\pm$0.2 97.5 $\pm$0.1 97.7 $\pm$0.1 93.8 $\pm$0.3 97.8 $\pm$0.1 96.7 $\pm$0.2 96.7 $\pm$0.1 97.6 $\pm$0.1 97.6 $\pm$0.1 95.3 $\pm$0.2 98.3 $\pm$0.1 97.8 $\pm$0.1 97.9 $\pm$0.1 96.6 $\pm$0.2 97.3 $\pm$0.1
C13 70.9 $\pm$0.8 73.7 $\pm$1.1 74.1 $\pm$1.0 70.7 $\pm$0.7 74.0 $\pm$0.8 72.6 $\pm$1.0 72.7 $\pm$1.0 73.0 $\pm$0.7 72.4 $\pm$1.2 72.3 $\pm$0.9 71.1 $\pm$1.1 75.0 $\pm$0.7 74.3 $\pm$0.6 74.4 $\pm$0.7 70.2 $\pm$0.8 74.1 $\pm$1.1
C14 83.5 $\pm$0.3 84.0 $\pm$0.5 82.4 $\pm$0.3 80.4 $\pm$0.3 81.9 $\pm$0.4 82.3 $\pm$0.4 81.9 $\pm$0.5 83.2 $\pm$0.4 80.1 $\pm$0.3 84.0 $\pm$0.3 84.8 $\pm$0.2 85.4 $\pm$0.3 81.4 $\pm$0.4 83.2 $\pm$0.4
Table 20: Classification performance per dataset (above 10K). Best score before rounding is bolded. We report average AUROC with 95% CIs. Models: CARTE (CRT), CatBoost (CTB), TabICL (ICL), LightGBM (LGB), RealMLP (MLP), TabDPT (DPT), TabM (TBM), TabPFNv2 (PFN), RandomForest (RF), TabSTAR (STR), XGBoost (XGB). Unlimited models are marked with a ’!’.
ID CRT CTB CTB! DPT ICL ICL! LGB LGB! MLP MLP! PFN STR STR! TBM TBM! XGB XGB!
C01 88.4 $\pm$0.3 90.3 $\pm$0.4 90.5 $\pm$0.4 90.2 $\pm$0.3 90.7 $\pm$0.4 90.9 $\pm$0.4 89.6 $\pm$0.4 89.8 $\pm$0.4 90.2 $\pm$0.3 90.6 $\pm$0.4 90.3 $\pm$0.3 90.8 $\pm$0.3 91.2 $\pm$0.3 90.5 $\pm$0.3 90.8 $\pm$0.4 90.2 $\pm$0.4 90.4 $\pm$0.3
C02 97.4 $\pm$0.4 98.3 $\pm$0.3 92.3 $\pm$0.6 95.2 $\pm$0.4 96.9 $\pm$0.5 97.3 $\pm$0.5 98.3 $\pm$0.3 96.6 $\pm$0.5 98.5 $\pm$0.2 97.9 $\pm$0.5 98.4 $\pm$0.2 97.2 $\pm$0.4 98.5 $\pm$0.3 97.6 $\pm$0.3 98.2 $\pm$0.3
C03 82.5 $\pm$0.3 81.2 $\pm$0.6 81.5 $\pm$0.8 85.4 $\pm$0.3 81.4 $\pm$0.3 81.2 $\pm$0.3 79.9 $\pm$0.3 80.4 $\pm$0.3 82.7 $\pm$0.3 83.4 $\pm$0.3 82.4 $\pm$0.3 83.0 $\pm$0.3 83.8 $\pm$0.5 83.8 $\pm$0.3 84.1 $\pm$0.4 82.1 $\pm$0.3 82.3 $\pm$0.3
C07 90.1 $\pm$0.4 90.6 $\pm$0.3 91.0 $\pm$0.4 89.9 $\pm$0.2 86.7 $\pm$0.3 87.1 $\pm$0.3 89.3 $\pm$0.4 90.0 $\pm$0.3 90.2 $\pm$0.2 90.5 $\pm$0.4 89.8 $\pm$0.4 91.5 $\pm$0.3 91.9 $\pm$0.3 91.1 $\pm$0.3 91.6 $\pm$0.3 89.3 $\pm$0.5 89.9 $\pm$0.5
C08 90.8 $\pm$1.6 93.2 $\pm$1.0 93.1 $\pm$1.2 93.2 $\pm$1.4 95.1 $\pm$0.7 95.3 $\pm$0.6 93.2 $\pm$0.9 93.4 $\pm$0.9 91.9 $\pm$1.3 91.9 $\pm$1.3 91.3 $\pm$1.1 93.5 $\pm$1.5 95.1 $\pm$0.9 92.1 $\pm$1.5 93.9 $\pm$1.1 94.4 $\pm$0.7 94.4 $\pm$0.9
C09 82.6 $\pm$1.4 85.9 $\pm$1.2 81.3 $\pm$1.2 84.3 $\pm$1.0 86.3 $\pm$1.1 82.9 $\pm$1.3 85.6 $\pm$1.2 81.3 $\pm$2.5 84.8 $\pm$0.9 82.5 $\pm$1.1 94.0 $\pm$0.8 96.2 $\pm$0.3 83.6 $\pm$1.0 86.1 $\pm$1.3 83.2 $\pm$1.2 85.5 $\pm$1.2
C11 78.4 $\pm$0.5 80.7 $\pm$0.4 81.8 $\pm$0.4 79.6 $\pm$0.4 81.5 $\pm$0.5 82.3 $\pm$0.4 79.2 $\pm$0.4 80.5 $\pm$0.5 79.6 $\pm$0.5 81.0 $\pm$0.3 81.4 $\pm$0.4 79.2 $\pm$0.5 81.0 $\pm$0.5 81.5 $\pm$0.3 82.8 $\pm$0.4 80.2 $\pm$0.5 81.4 $\pm$0.4
C12 96.0 $\pm$0.2 97.7 $\pm$0.1 98.5 $\pm$0.1 93.8 $\pm$0.3 97.8 $\pm$0.1 98.6 $\pm$0.1 96.7 $\pm$0.1 79.7 $\pm$1.1 97.6 $\pm$0.1 98.5 $\pm$0.1 98.3 $\pm$0.1 99.1 $\pm$0.1 97.9 $\pm$0.1 98.7 $\pm$0.1 97.3 $\pm$0.1 98.4 $\pm$0.1
C13 70.9 $\pm$0.8 74.1 $\pm$1.0 76.7 $\pm$0.8 70.7 $\pm$0.7 74.0 $\pm$0.8 76.7 $\pm$0.8 72.7 $\pm$1.0 75.8 $\pm$0.9 72.4 $\pm$1.2 76.4 $\pm$0.6 72.3 $\pm$0.9 75.0 $\pm$0.7 77.9 $\pm$1.0 74.4 $\pm$0.7 77.6 $\pm$1.0 74.1 $\pm$1.1 76.9 $\pm$0.9
C14 84.0 $\pm$0.5 84.8 $\pm$0.4 82.4 $\pm$0.3 80.4 $\pm$0.3 81.3 $\pm$0.3 82.3 $\pm$0.4 83.5 $\pm$0.3 83.2 $\pm$0.4 84.3 $\pm$0.3 84.0 $\pm$0.3 85.1 $\pm$0.3 85.4 $\pm$0.3 86.2 $\pm$0.3 83.2 $\pm$0.4 84.2 $\pm$0.4
Table 21: Regression performance per dataset (up to 10K). Best score before rounding is bolded. We report average $R^{2}$ with 95% CIs. Models: CARTE (CRT), CatBoost (CTB), LightGBM (LGB), RealMLP (MLP), TabM (TBM), TabPFNv2 (PFN), RandomForest (RF), TabSTAR (STR), XGBoost (XGB). Tuned models are marked with a ’+’.
ID CRT CTB CTB+ LGB LGB+ MLP MLP+ PFN RF STR TBM TBM+ XGB XGB+
R01 100.0 $\pm$0.0 100.0 $\pm$0.0 100.0 $\pm$0.0 100.0 $\pm$0.0 100.0 $\pm$0.0 99.9 $\pm$0.1 100.0 $\pm$0.0 100.0 $\pm$0.0 100.0 $\pm$0.0 99.8 $\pm$0.1 99.8 $\pm$0.1 99.9 $\pm$0.1 100.0 $\pm$0.0
R02 98.3 $\pm$1.0 98.3 $\pm$0.9 98.1 $\pm$1.0 98.1 $\pm$1.0 98.2 $\pm$1.0 98.2 $\pm$1.0 98.4 $\pm$0.9 98.0 $\pm$1.0 98.0 $\pm$1.1 99.0 $\pm$0.8 99.0 $\pm$0.8 97.9 $\pm$1.1 98.3 $\pm$0.9
R03 71.8 $\pm$0.5 73.8 $\pm$0.4 74.1 $\pm$0.3 72.2 $\pm$0.4 72.4 $\pm$0.4 69.7 $\pm$0.7 70.1 $\pm$0.8 74.9 $\pm$0.4 68.4 $\pm$0.6 70.9 $\pm$0.8 75.3 $\pm$0.5 75.4 $\pm$0.5 69.1 $\pm$0.5 73.7 $\pm$0.5
R04 86.2 $\pm$7.4 85.1 $\pm$7.1 85.7 $\pm$6.5 84.6 $\pm$6.2 88.7 $\pm$4.4 90.3 $\pm$3.2 86.1 $\pm$5.2 85.1 $\pm$6.8 81.5 $\pm$8.7 84.0 $\pm$3.1 84.1 $\pm$3.1 85.8 $\pm$7.5 87.5 $\pm$3.7
R05 93.2 $\pm$0.8 89.5 $\pm$1.3 90.0 $\pm$1.2 87.4 $\pm$1.5 88.2 $\pm$1.4 90.9 $\pm$1.2 91.2 $\pm$1.5 92.7 $\pm$0.9 86.7 $\pm$1.3 93.2 $\pm$1.2 92.1 $\pm$1.1 92.3 $\pm$1.1 87.1 $\pm$1.4 90.1 $\pm$1.3
R06 97.7 $\pm$0.8 98.1 $\pm$0.5 98.2 $\pm$0.5 97.7 $\pm$0.7 97.8 $\pm$0.7 97.5 $\pm$0.9 97.8 $\pm$0.9 98.5 $\pm$0.6 97.2 $\pm$1.0 98.1 $\pm$0.3 97.8 $\pm$0.8 97.9 $\pm$0.8 97.1 $\pm$0.7 97.9 $\pm$0.7
R07 71.6 $\pm$1.3 66.9 $\pm$0.8 67.8 $\pm$0.8 62.4 $\pm$1.0 63.4 $\pm$1.0 68.9 $\pm$0.8 67.4 $\pm$1.3 62.2 $\pm$1.3 61.5 $\pm$1.2 71.1 $\pm$1.8 67.9 $\pm$0.7 69.6 $\pm$0.9 61.9 $\pm$1.5 68.8 $\pm$0.8
R08 93.2 $\pm$0.8 93.0 $\pm$0.7 93.2 $\pm$0.7 92.9 $\pm$0.7 93.0 $\pm$0.7 92.9 $\pm$0.8 92.7 $\pm$0.8 93.9 $\pm$0.7 92.2 $\pm$0.7 92.8 $\pm$0.6 93.4 $\pm$0.7 93.3 $\pm$0.7 92.3 $\pm$0.7 93.1 $\pm$0.7
R09 89.2 $\pm$0.5 89.2 $\pm$0.6 89.3 $\pm$0.4 88.7 $\pm$0.5 88.9 $\pm$0.5 88.6 $\pm$0.4 89.2 $\pm$0.5 89.8 $\pm$0.5 88.6 $\pm$0.7 88.8 $\pm$0.5 89.5 $\pm$0.7 89.7 $\pm$0.7 88.2 $\pm$0.7 89.3 $\pm$0.5
R10 99.5 $\pm$0.2 99.0 $\pm$0.6 98.9 $\pm$0.6 98.5 $\pm$0.3 98.5 $\pm$0.3 99.3 $\pm$0.2 99.4 $\pm$0.2 99.1 $\pm$0.3 98.1 $\pm$0.3 99.5 $\pm$0.1 98.9 $\pm$0.7 99.1 $\pm$0.4 98.6 $\pm$0.3 99.4 $\pm$0.2
R11 94.2 $\pm$0.9 94.0 $\pm$0.9 93.6 $\pm$0.9 93.6 $\pm$1.0 94.2 $\pm$1.0 94.4 $\pm$1.0 94.5 $\pm$0.9 92.5 $\pm$1.3 93.9 $\pm$1.1 94.5 $\pm$0.8 94.6 $\pm$0.8 93.3 $\pm$1.1 94.2 $\pm$0.8
R12 98.5 $\pm$0.3 98.5 $\pm$0.2 98.3 $\pm$0.3 98.4 $\pm$0.2 98.4 $\pm$0.3 98.4 $\pm$0.2 98.5 $\pm$0.2 98.0 $\pm$0.3 98.3 $\pm$0.2 98.6 $\pm$0.2 98.7 $\pm$0.2 98.2 $\pm$0.3 98.5 $\pm$0.2
R13 52.8 $\pm$2.1 57.6 $\pm$2.7 58.2 $\pm$2.6 54.7 $\pm$2.8 55.0 $\pm$2.6 55.4 $\pm$2.5 55.1 $\pm$2.5 53.8 $\pm$2.7 51.8 $\pm$3.5 53.4 $\pm$3.0 58.7 $\pm$1.9 58.4 $\pm$2.1 49.0 $\pm$3.3 57.5 $\pm$2.7
R14 96.6 $\pm$0.2 97.3 $\pm$0.1 97.3 $\pm$0.1 97.0 $\pm$0.1 97.0 $\pm$0.1 96.8 $\pm$0.1 96.7 $\pm$0.2 96.2 $\pm$0.2 96.8 $\pm$0.1 96.4 $\pm$0.1 96.3 $\pm$0.1 96.3 $\pm$0.1 97.1 $\pm$0.1 97.1 $\pm$0.1
R15 79.5 $\pm$0.9 79.9 $\pm$1.1 77.8 $\pm$0.9 77.7 $\pm$0.9 76.2 $\pm$1.4 77.4 $\pm$1.5 79.4 $\pm$1.0 77.3 $\pm$1.2 78.9 $\pm$1.5 80.6 $\pm$0.7 79.9 $\pm$0.7 76.8 $\pm$1.2 80.3 $\pm$1.0
R16 98.7 $\pm$0.0 98.7 $\pm$0.0 98.7 $\pm$0.0 98.7 $\pm$0.0 98.7 $\pm$0.0 97.6 $\pm$0.1 97.8 $\pm$0.1 98.8 $\pm$0.0 97.7 $\pm$0.1 98.7 $\pm$0.0 97.9 $\pm$0.1 97.9 $\pm$0.1 97.7 $\pm$0.1 97.8 $\pm$0.1
R17 95.3 $\pm$2.0 94.8 $\pm$2.4 95.2 $\pm$2.0 94.6 $\pm$2.3 95.2 $\pm$2.0 94.4 $\pm$2.4 95.4 $\pm$1.9 95.1 $\pm$2.1 94.2 $\pm$2.5 94.7 $\pm$2.3 96.7 $\pm$1.1 96.7 $\pm$1.1 94.5 $\pm$2.6 95.4 $\pm$1.9
R18 98.1 $\pm$0.4 98.2 $\pm$0.4 98.2 $\pm$0.4 98.1 $\pm$0.4 98.1 $\pm$0.4 98.0 $\pm$0.5 97.7 $\pm$0.5 93.1 $\pm$1.3 97.7 $\pm$0.4 97.4 $\pm$0.6 98.2 $\pm$0.3 98.1 $\pm$0.4 97.5 $\pm$0.8 98.3 $\pm$0.4
R19 85.5 $\pm$0.8 85.8 $\pm$0.9 85.2 $\pm$0.9 85.2 $\pm$0.9 82.7 $\pm$0.8 83.3 $\pm$1.3 84.0 $\pm$1.1 85.3 $\pm$0.9 83.8 $\pm$1.1 85.0 $\pm$1.0 85.1 $\pm$0.9 83.9 $\pm$1.0 82.4 $\pm$2.0
R20 96.4 $\pm$0.5 96.0 $\pm$0.6 96.2 $\pm$0.5 95.7 $\pm$0.5 95.8 $\pm$0.5 95.3 $\pm$0.8 96.0 $\pm$0.5 93.9 $\pm$1.0 95.7 $\pm$0.6 95.6 $\pm$0.9 95.9 $\pm$0.7 95.8 $\pm$0.7 95.3 $\pm$0.9 96.2 $\pm$0.5
R21 92.3 $\pm$1.2 92.5 $\pm$1.1 92.6 $\pm$1.1 92.3 $\pm$1.2 92.3 $\pm$1.2 92.0 $\pm$1.1 91.7 $\pm$1.3 93.3 $\pm$1.1 91.8 $\pm$1.1 92.3 $\pm$1.0 91.6 $\pm$1.4 92.1 $\pm$1.2 91.6 $\pm$1.0 92.6 $\pm$1.0
R22 45.2 $\pm$5.5 45.2 $\pm$5.9 42.4 $\pm$5.6 44.7 $\pm$5.5 43.6 $\pm$6.4 44.4 $\pm$6.3 43.8 $\pm$5.5 39.1 $\pm$6.4 39.7 $\pm$5.5 47.9 $\pm$5.5 48.1 $\pm$5.4 36.1 $\pm$7.1 46.8 $\pm$5.8
R23 85.0 $\pm$1.8 85.4 $\pm$1.2 85.5 $\pm$1.1 84.3 $\pm$1.5 85.5 $\pm$1.2 83.3 $\pm$1.8 85.6 $\pm$1.5 84.8 $\pm$1.6 81.8 $\pm$1.5 83.0 $\pm$1.6 86.4 $\pm$0.9 86.6 $\pm$0.8 83.3 $\pm$1.2 85.1 $\pm$1.2
R24 86.0 $\pm$0.6 84.0 $\pm$0.8 85.8 $\pm$0.7 78.8 $\pm$0.9 79.5 $\pm$0.9 84.6 $\pm$0.8 82.8 $\pm$0.9 70.3 $\pm$1.6 79.7 $\pm$1.0 81.8 $\pm$1.8 84.9 $\pm$0.9 85.5 $\pm$0.8 82.9 $\pm$0.7 85.9 $\pm$0.9
R25 94.3 $\pm$0.6 95.4 $\pm$0.5 95.3 $\pm$0.5 95.2 $\pm$0.6 95.2 $\pm$0.5 93.2 $\pm$0.6 94.1 $\pm$0.7 86.7 $\pm$1.0 94.8 $\pm$0.5 94.8 $\pm$0.5 94.3 $\pm$0.4 94.5 $\pm$0.4 94.6 $\pm$0.6 95.4 $\pm$0.5
R26 99.8 $\pm$0.1 99.4 $\pm$0.1 99.4 $\pm$0.1 99.2 $\pm$0.2 99.3 $\pm$0.2 99.9 $\pm$0.1 99.9 $\pm$0.0 99.8 $\pm$0.1 99.0 $\pm$0.2 99.6 $\pm$0.1 98.1 $\pm$0.7 98.7 $\pm$0.8 99.1 $\pm$0.2 99.5 $\pm$0.1
R27 81.9 $\pm$1.6 85.2 $\pm$1.3 85.3 $\pm$1.4 84.4 $\pm$1.4 84.8 $\pm$1.3 83.3 $\pm$1.5 76.7 $\pm$11.2 82.2 $\pm$1.6 84.8 $\pm$1.3 82.0 $\pm$1.6 83.8 $\pm$1.5 84.0 $\pm$1.7 83.2 $\pm$1.2 85.5 $\pm$1.2
[…]